# Mapping Cone Connections and their Yang-Mills Functional

### v1

@@ 2.4 Uniqueness of cone Yang-Mills solutions for a fixed connection
In general, given an arbitrary condition connection $A$, there does not exist a $B$ that would make $(A,B)$ a critical point of the cone Yang-Mills functional. To illustrate this in a simple example, let $M$ be a closed manifold and $\zeta=0\,$. Then by Lemma 2.2, we know that $A$ must be a Yang-Mills connection and $d_{A}B=0\,$. Hence, there exists no $B$ that would give a cone Yang-Mills solution if $A$ is not Yang-Mills.
However, it is interesting to ask that given a cone Yang-Mills solution $(A,B)$, how many different $B$’s with $A$ fixed would also be a cone Yang-Mills solution? In the case of $\zeta=0$ and $M$ closed, if $(A,B)$ is a critical point, then so are all $(A,B+\Phi)$ with $d_{A}\Phi=0$. For $\zeta\neq 0$, we are able to obtain results in two special cases. First, for two dimensions and suppose $A$ is a Yang-Mills connection, then we have the following:
@@
Suppose manifold $M$ is a closed Riemann surface and $\zeta$ is its volume form. For each Yang-Mills connection $A$, there exists a unique $B$ such that $(A,B)$ is a critical point of the cone Yang-Mills functional. Actually, such $(A,B)$ is always a zero point.
@@
Since $A$ is Yang-Mills, $F_{A}=\zeta\Phi$ with $d_{A}\Phi=0$. So we can take $B=-\Phi$ then $(A,B)$ is a zero point of the cone Yang-Mills functional.
Now suppose $(A,B)$ is a critical point. As $\zeta$ is a volume form, the condition $\zeta^{*}(F_{A}+\zeta B)+d_{A}^{*}d_{A}B=0$ becomes
$\Phi+B+d_{A}^{*}d_{A}B$ $=0$ (2.23)
Let $d_{A}$ act on both sides, we have $d_{A}B=-d_{A}d_{A}^{*}d_{A}B$. So
@@
So $\langle d_{A}B,d_{A}^{*}d_{A}B\rangle$ has to vanish. It follows that $\|d_{A}B\|^{2}=0$, then $d_{A}B=0$, and by (2.23), $B=-\Phi$. So such $B$ is unique when $A$ is Yang-Mills. ∎
@@
Suppose $A$ is a connection on the associated adjoint bundle $Ad\,P$ over a symplectic manifold $(M^{2n},\omega)$ and the Riemannian metric is compatible with $\omega$. Take $\zeta=\omega$. Then there is at most one $B$ that satisfies $[B,d_{A}B]=0$ and such that $(A,B)$ is a critical point of the cone Yang-Mills functional. In particular, if the structure group is abelian, there is at most one critical point $(A,B)$ for any given connection $A$.
@@
Assume $\dim M=2n$. With respect to a compatible metric, $*\zeta=*\omega=\frac{1}{(n-1)!}\omega^{n-1}$, and $\zeta^{*}\zeta=n$
Suppose both $(A,B_{1})$ and $(A,B_{2})$ are critical points of the cone Yang-Mills functional, and $[B_{1},d_{A}B_{1}]=[B_{2},d_{A}B_{2}]=0$. Then (2.9)-(2.10) imply
@@
This case avoids considering the term $[B,d_{A}B]$. If $[B,d_{A}B]=0$ for all critical points $(A,B)$, then we would have obtained the uniqueness of $B$ when $\zeta$ is the symplectic structure. But as we shall see in Example 3.2 in the next section, a cone Yang-Mills solution in general need not satisfy $[B,d_{A}B]=0$.
@@ 3.1 Two-dimensional solutions with abelian gauge group
Thus, $\zeta^{*}\!\left(F_{A}+\zeta B\right)=cc^{\prime}+c\,d^{*}df$. Plugging this into (3.2), we find that $cc^{\prime}$ is a $d^{*}$-exact constant; hence, $cc^{\prime}$ must be zero.
@@
Suppose $M$ is a closed Riemann surface, $\zeta$ is a non-exact two-form on $M$, and the structure group is abelian. Then the only critical value of the cone Yang-Mills functional is zero.
Now we consider the case $c^{\prime}=0$, that is, $\zeta=-d*df$ is a $d$-exact form. Then $d^{*}d(cf+B)=0$ and so $cf+B$ is a constant. We can thus write $B=-cf+c^{\prime\prime}$ for some constant $c^{\prime\prime}$. Hence, by (3.3), $F_{A}=c\,\omega-\zeta B=c\,\omega+(c^{\prime\prime}-cf)d*df$. Therefore, we find that the constants $c$ and $c^{\prime\prime}$ parametrize the critical points. And finally, in the special case where $f=0$, i.e. $\zeta$ vanishes, the critical points must satisfy $F_{A}=c\,\omega$ and $B=c^{\prime\prime}$.
@@ 3.2 Three-dimensional solutions from duality relations
Recall in four dimensions, there are special Yang-Mills solutions that satisfy the first-order self-dual/anti-self-dual conditions, $*\,F_{A}=\pm F_{A}\,$. The intuition that cone Yang-Mills functional can be interpreted as a dimensional reduction of the Yang-Mills functional suggests an analogous duality condition $*_{\mathcal{C}}\,\mathcal{F}=\pm\mathcal{F}$ over 3-manifolds, where $\mathcal{F}=(F_{A}+\zeta B)-\theta d_{A}B$. With the $*_{\mathcal{C}}$ acting on $\operatorname{Cone}^{k}(\zeta)$ given by (2.1), we have
$*_{\mathcal{C}}\,\mathcal{F}$ $=*_{\mathcal{C}}\left[(F_{A}+\zeta B)-\theta d_{A}B\right]$
@@
Hence, we find that $*_{\mathcal{C}}\,\mathcal{F}=\mp\mathcal{F}$ implies
@@
$\|\mathcal{F}\|_{\mathcal{C}}^{2}$ $=\int_{M}c_{G}\,\mathrm{tr}[(F_{A}+\zeta B)\wedge*(F_{A}+\zeta B)\,+\,d_{A}B\wedge*d_{A}B)]$
@@
which is the action of a two-dimensional topologcial $BF$-type theory. We note that any infinitesimal local variation of $(A,B)$ away from the boundary does not affect the bound. Hence, $F_{A}+\zeta B=\pm*d_{A}B$ must be a critical point of the cone Yang-Mills functional.
@@
The Taub-NUT gravitational instanton solution can be thought of as a self-dual Yang-Mills solutions of a tangent bundle with structure group $SU(2)\subset SO(4)$. With a point removed, the four-dimensional space can be considered as a circle bundle, $S^{1}\to X\to\mathbb{R}^{3}-\{0\}$ (see, for example, the description in [LeBrun]). Dimensionally reduce this solution leads to a non-abelian cone Yang-Mills solution on $M=\mathbb{R}^{3}-\{0\}$ satisfying (3.4).
@@
These result in the following connection 1-forms (for a reference, see, [Oh])
@@
We can consider dimensionally reducing the three-dimensional duality condition $*_{\mathcal{C}}\,\mathcal{F}=\mp\mathcal{F}$ over another circle down to a two-dimensional manifold $N$. For the metric, we will assume $ds_{M}^{2}=ds_{N}^{2}+\chi^{2}$ where $\chi$ is the global connection one-form (i.e. the global angular form) of the circle bundle $M$ over $N$. The dimensionally reduced three-dimensional local connection can then be expressed as $A+\chi b$ where $A\in\Omega^{1}(N,Ad\,P)$ and $b\in\Omega^{0}(N,Ad\,P)$. The three-dimensional cone curvature then takes the form
$\mathcal{F}=F_{A+\chi b}+\zeta B-\theta d_{A+\chi b}B$ $=F_{A}+\zeta B+(d\chi)b-\chi d_{A}b-\theta d_{A}B-\theta\chi[b,B]\,,$
@@
$*_{\mathcal{C}}\mathcal{F}=\theta\chi\left[*\left(F_{A}+\zeta B+(d\chi)b\right)\right]-\theta*d_{A}b+\chi*d_{A}B-*[b,B]\,.$
The condition $*_{\mathcal{C}}\mathcal{F}=\mp\mathcal{F}$ then implies the following two equations:
$F_{A}+\zeta B+(d\chi)b$ =±$=\pm*[b,B]$ (3.15)
$*\,d_{A}B$ =±$=\pm\,d_{A}b$ (3.16)
@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
Now if $\zeta=0$, the cone-flat condition reduces to (1) $F_{A}=0$, i.e. $A$ is a flat connection, and (2) $d_{A}B=0$, that is, $B$ is a covariantly constant section. Since we can always set $B=0$, the $\zeta=0$ cone-flat bundles are just the usual flat bundles, and flat bundles are well-known to be classified by the conjugacy classes of homomorphisms from $\pi_{1}(M)$ to the structure group $G$ of the fiber bundle (see, for example [Morita]). As the classification for the $\zeta=0$ case is well-known, we will in the following always assume that $\zeta\neq 0$.
To simplify our consideration, we will additionally assume that $\zeta$ is a non-degenerate two-form. Being $d$-closed and non-degenerate, $\zeta$ is then a symplectic structure and $M$ must be even-dimensional. It is worthwhile to note that when $\dim M\geq 4$, the first cone-flat equation in (4.1) together with the Bianchi identity imply
@@
If $\zeta$ is non-degenerate, then (4.2) automatically implies the second cone-flat equation $d_{A}B=0\,$. Hence, the $\zeta$ non-degenerate assumption will simplify our consideration and allow us to only consider the first condition, $F_{A}=-\zeta\,B$, i.e. that the curvature is proportional to $\zeta$, if $\dim M\geq 4$.
When $M$ is a two-dimensional Riemann surface, $\zeta$ being non-degenerate implies that it is the product of a nowhere vanishing function times the volume form. If we take $\zeta$ to be volume form, then we saw in (1.12) that the cone-flat condition becomes equivalent to the Yang-Mills equation for $A$ when $g$ is a compatible metric, and $B=-*\,F_{A}$. For two-dimensional Yang-Mills connections, Atiyah-Bott [Atiyah-Bott] gave a classification of principal bundles that carry a Yang-Mills connection. So we ask whether Atiyah-Bott’s classification can be extended to cone-flat connections with respect to a closed, non-degenerate two-form $\zeta$ for $\dim M\geq 2$. Indeed, we obtain the following result for principal bundles which can be thought of as a generalization of Atiyah-Bott’s result on Riemann surfaces to higher even-dimensional spaces.
@@
We will provide a more explicit description of $\Gamma$ in terms of path spaces. To do so, we first introduce some notation.
@@
which is a one-parameter family of subpaths, parametrized by $t$, all with the same starting point $\alpha(0)$, and ending point being at $\alpha(t)$. The following subsets of $\Psi(M)$ will be of interest:
@@
    $\Psi^{a}_{a}(M)/\Psi^{a}_{a,0}(M)=\pi_{1}(M)$ is the fundamental group where the identity element is the equivalence class of the constant path at $a$, and the inverse reverses the path orientation, i.e. $\alpha^{-1}(t)=\alpha(1-t)\,$.
@@
    $\Psi^{a}_{a,0}(M)/\Psi^{a}_{a,\zeta}(M)=\mathbb{R}/H$. As $\zeta\neq 0$, for any real number $c$ there exists some $D$ such that $\int_{D}\zeta=c$. So $\Psi^{a}_{a,0}(M)/\Psi^{a}_{a,\zeta}(M)$ is equivalent to a quotient group $\mathbb{R}/H$ by the identification of a loop $\gamma\mapsto[\,\int_{D}\zeta\,]$, where
@@
    $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)=:\Gamma$. This is the extension of $\pi_{1}(M)$ by $\mathbb{R}/\overline{H}$. In the case where $H$ is dense and thus $\mathbb{R}/H$ is not a Lie group, $\Gamma=\pi_{1}(M)$.
@@
Case 1. When $H^{+}$ is empty, $H=0$ and $\mathbb{R}/H=\mathbb{R}$. In this case $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ is an $\mathbb{R}$-extension of $\pi_{1}(M)$. For Riemann surfaces, this occurs when the genus $g\geq 1$.
Case 2. When $H^{+}$ has a minimal number, $H\simeq\mathbb{Z}$ and $\mathbb{R}/H\simeq S^{1}$. In this case $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ is an $S^{1}$-extension of $\pi_{1}(M)$. For Riemann surfaces, this occurs when the genus $g=0$ and $\zeta$ is not an exact form.
@@
In the remainder of this subsection, we prove the classification theorem. Our proof will be similar to that given by Morrison [Morrison] for Yang-Mills bundles in the two-dimensional case.
We will first consider the case that $\mathbb{R}/H$ is a Lie group. By the discussion above, $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ is either an $\mathbb{R}$ or an $S^{1}$ extension of $\pi_{1}(M)$.
@@
Step 1. Given a homomorphism $\rho:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G$, construct a corresponding principal $G$-bundle $P_{\rho}$ with a connection $A_{\rho}$ and a covariantly constant section $B_{\rho}\in\Omega^{0}(M,Ad\,P)$.
Let $\Psi_{a}(M)$ be the space of classes of paths starting at $a$ on $M$. Two paths $\delta_{1},\delta_{2}$ are identified if $\delta_{1}(1)=\delta_{2}(1)$ and $(\delta_{2})^{-1}\delta_{1}\in\Psi^{a}_{a,\zeta}(M)$. This $\Psi_{a}(M)$ is a principal bundle over $M$. Its structure group is $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$, and the projection is $\tau:\Psi_{a}(M)\to M,\delta\mapsto\delta(1)$. An element $\chi\in\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ acts on $\delta\in\Psi_{a}(M)$ on the right by $\delta\mapsto\delta\chi$.
Given a homomorphism $\rho:\Gamma\to G$, we define a principal bundle as a $G$-bundle associated to $\Psi_{a}(M)$:

### v2

@@ 2.4 Uniqueness of cone Yang-Mills solutions for a fixed connection
In general, given an arbitrary connection $A$, there does not exist a $B$ that would make $(A,B)$ a cone Yang-Mills solution. To illustrate this in a simple example, let $M$ be a closed manifold and $\zeta=0\,$. Then by Lemma 2.2, we know that $A$ must be a Yang-Mills connection and $d_{A}B=0\,$. Hence, there exists no $B$ that would give a cone Yang-Mills solution if $A$ is not Yang-Mills.
However, it is interesting to ask that given a cone Yang-Mills solution $(A,B)$, how many different $B$’s with $A$ fixed would also be a cone Yang-Mills solution? In the case of $\zeta=0$ and $M$ closed, if $(A,B)$ is a cone Yang-Mills solution, then so are all $(A,B+\Phi)$ with $d_{A}\Phi=0$. For $\zeta\neq 0$, we are able to obtain results in two special cases. First, for two dimensions and suppose $A$ is a Yang-Mills connection, then we have the following:
@@
Suppose $M$ is a closed Riemann surface and $\zeta$ is its volume form. For each Yang-Mills connection $A$, there exists a unique $B$ such that $(A,B)$ is a cone Yang-Mills solution. In fact, such a pair $(A,B)$ is always cone-flat.
@@
Since $A$ is a Yang-Mills connection on a Riemann surface, $F_{A}=\zeta\Phi$ with $d_{A}\Phi=0$. So we can choose $B=-\Phi$ and then $(A,B)$ is a cone-flat solution. We will show below that this is the unique cone Yang-Mills solution for a fixed $A$, a Yang-Mills connection.
Generally, suppose $(A,B)$ is a cone Yang-Mills solution. As $\zeta$ is a volume form, the condition $\zeta^{*}(F_{A}+\zeta B)+d_{A}^{*}d_{A}B=0$ becomes
$\Phi+B+d_{A}^{*}d_{A}B=0\,.$ (2.23)
Let $d_{A}$ act on both sides. This implies $d_{A}B=-d_{A}d_{A}^{*}d_{A}B$, and therefore,
@@
So $\langle d_{A}B,d_{A}d_{A}^{*}d_{A}B\rangle$ has to vanish. It follows that $\|d_{A}B\|^{2}=0$, which implies, $d_{A}B=0$, and by (2.23), $B=-\Phi$. So such $B$ is unique when $A$ is Yang-Mills. ∎
@@
Suppose $A$ is a connection on the associated adjoint bundle $Ad\,P$ over a symplectic manifold $(M^{2n},\omega)$ and the Riemannian metric is compatible with $\omega$. Take $\zeta=\omega$. Then there is at most one $B$ that satisfies $[B,d_{A}B]=0$ and such that $(A,B)$ is a cone Yang-Mills solution. In particular, if the structure group is abelian, there is at most one cone Yang-Mills solution pair $(A,B)$ for any given connection $A$.
@@
Let $\dim M=2n$. With respect to a compatible metric, $*\,\zeta=*\,\omega=\frac{1}{(n-1)!}\omega^{n-1}$, and $\zeta^{*}\zeta=n$
Suppose both $(A,B_{1})$ and $(A,B_{2})$ are cone Yang-Mills solutions. Assume also $[B_{1},d_{A}B_{1}]=[B_{2},d_{A}B_{2}]=0$ which is identically true when the structure group is abelian. Then (2.9)-(2.10) imply for $B_{1}$,
@@
The above proposition avoids considering the term $[B,d_{A}B]$. If $[B,d_{A}B]=0$ holds true for all cone Yang-Mills solutions $(A,B)$, then we would have obtained the uniqueness of $B$ when $\zeta$ is the symplectic structure. But as we shall see in Example 3.2 in the next section, a cone Yang-Mills solution in general need not satisfy $[B,d_{A}B]=0$.
@@ 3.1 Two-dimensional solutions with abelian gauge group
and therefore,
$\zeta^{*}\!\left(F_{A}+\zeta B\right)=cc^{\prime}+c\,d^{*}df\,.$ (3.4)
Plugging this into (3.2), we find that $cc^{\prime}$ is a $d^{*}$-exact constant; hence, $cc^{\prime}$ must be zero.
@@
Suppose $M$ is a closed Riemann surface, $\zeta$ is a non-exact two-form on $M$, and the structure group is abelian. Then all cone Yang-Mills solutions are cone-flat.
Now we consider the case $c^{\prime}=0$, that is, $\zeta=-d*df$ is a $d$-exact form. Together, (3.2) and (3.4) imply $d^{*}d(cf+B)=0$, and so $cf+B$ is a constant. We can thus write $B=-cf+c^{\prime\prime}$ for some constant $c^{\prime\prime}$. Hence, by (3.3), $F_{A}=c\,\omega-\zeta B=c\,\omega+(c^{\prime\prime}-cf)d*df$. Therefore, we find that the constants $c$ and $c^{\prime\prime}$ parametrize the cone Yang-Mills solutions. And finally, in the special case where $f=0$, i.e. $\zeta$ vanishes, the critical points must satisfy $F_{A}=c\,\omega$ and $B=c^{\prime\prime}$.
@@ 3.2 Three-dimensional solutions from duality relations
Recall in four dimensions, there are special Yang-Mills solutions that satisfy the first-order self-dual/anti-self-dual conditions, $*\,F_{A}=\pm\,F_{A}\,$. The intuition that cone Yang-Mills functional can be interpreted as a dimensional reduction of the Yang-Mills functional suggests an analogous duality condition $*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}=\pm\,\mathcal{F}_{\mathcal{A}}$ over 3-manifolds, where $\mathcal{F}_{\mathcal{A}}=(F_{A}+\zeta B)-\theta d_{A}B$. With the $*_{\mathcal{C}}$ acting on $\operatorname{Cone}^{k}(\zeta)$ given by (2.1), we have
$*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}$ $=*_{\mathcal{C}}\left[(F_{A}+\zeta B)-\theta d_{A}B\right]$
@@
Hence, we find that $*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}=\mp\,\mathcal{F}_{\mathcal{A}}$ implies
@@
$\|\mathcal{F}_{\mathcal{A}}\|_{\mathcal{C}}^{2}$ $=\int_{M}c_{G}\,\mathrm{tr}[(F_{A}+\zeta B)\wedge*(F_{A}+\zeta B)\,+\,d_{A}B\wedge*\,d_{A}B)]$
@@
which is the action of a two-dimensional topological $BF$-type theory. We note that any infinitesimal local variation of $(A,B)$ away from the boundary does not affect the bound. Hence, $F_{A}+\zeta B=\pm*d_{A}B$ must be a critical point of the cone Yang-Mills functional.
@@
The Taub-NUT gravitational instanton solution can be thought of as a self-dual Yang-Mills solutions of a tangent bundle with structure group $SU(2)\subset SO(4)$. With a point removed, the four-dimensional space can be considered as a circle bundle, $S^{1}\to X\to\mathbb{R}^{3}-\{0\}$ (see, for example, the description in [LeBrun]). Dimensionally reducing this solution leads to a non-abelian cone Yang-Mills solution on $M=\mathbb{R}^{3}-\{0\}$ satisfying (3.5).
@@
These result in the following connection 1-forms (for a reference, see [Oh])
@@
We can consider dimensionally reducing the three-dimensional duality condition $*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}=\mp\,\mathcal{F}_{\mathcal{A}}$ over another circle down to a two-dimensional manifold $N$. For the metric, we will assume $ds_{M}^{2}=ds_{N}^{2}+\chi^{2}$ where $\chi$ is the global connection one-form (i.e. the global angular form) of the circle bundle $M$ over $N$. The dimensionally reduced three-dimensional local connection can then be expressed as $A+\chi b$ where $A\in\Omega^{1}(N,Ad\,P)$ and $b\in\Omega^{0}(N,Ad\,P)$. The three-dimensional cone curvature then takes the form
$\mathcal{F}_{\mathcal{A}}=F_{A+\chi b}+\zeta B-\theta d_{A+\chi b}B$ $=F_{A}+\zeta B+(d\chi)b-\chi d_{A}b-\theta d_{A}B-\theta\chi[b,B]\,,$
@@
$*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}=\theta\chi\left[*\left(F_{A}+\zeta B+(d\chi)b\right)\right]-\theta*d_{A}b+\chi*d_{A}B-*[b,B]\,.$
The condition $*_{\mathcal{C}}\,\mathcal{F}_{\mathcal{A}}=\mp\,\mathcal{F}_{\mathcal{A}}$ then implies the following two equations:
$F_{A}+\zeta B+(d\chi)b$ =±$=\pm*[b,B]\,,$ (3.16)
$*\,d_{A}B$ =±$=\pm\,d_{A}b\,.$ (3.17)
@@ 4 Classification of cone-flat bundles with respect to a non-degenerate two-form
Now if $\zeta=0$, the cone-flat condition reduces to (1) $F_{A}=0$, i.e. $A$ is a flat connection, and (2) $d_{A}B=0$, that is, $B$ is a covariantly constant section. Since we can always set $B=0$, the $\zeta=0$ cone-flat bundles are just the usual flat bundles, and flat bundles are well-known to be classified by the conjugacy classes of homomorphisms from $\pi_{1}(M)$ to the structure group $G$ of the fiber bundle (see, for example [Morita]). As the classification for the $\zeta=0$ case is well-understood, we will in the following always assume that $\zeta\neq 0$.
To simplify our consideration, we will additionally assume that $\zeta$ is a non-degenerate two-form. Being $d$-closed and non-degenerate, $\zeta$ is then a symplectic structure and $M$ must be even-dimensional. Additionally, $\zeta$ being non-degenerate leads to two simplifications in considering cone-flat pairs $(A,B)$. First, non-degeneracy ensures that $\zeta$ nowhere vanishes, and hence, the first cone-flat equation $F_{A}=-\zeta B$, implies $B$ is determined by the connection $A$. Second, the second cone-flat equation $d_{A}B=0$ is automatically satisfied and redundant when $\dim M\geq 4$. For the first cone-flat equation together with the Bianchi identity imply
@@
If $\zeta$ is non-degenerate and $\dim M\geq 4$, then (4.2) implies $d_{A}B=0\,$. In all, $\zeta$ is non-degenerate simplifies our consideration of cone-flat pairs $(A,B)$ to checking only that that the curvature is proportional to $\zeta$, i.e. $F_{A}=-\zeta\,B$, and only if $\dim M=2$, we need to also check that the resulting $B$ is covariantly constant.
When $M$ is a two-dimensional Riemann surface, $\zeta$ being non-degenerate implies that it is the product of a nowhere vanishing function times the volume form. If we take $\zeta$ to be the volume form, then we saw in (1.12) that the cone-flat condition (4.1) becomes equivalent to the Yang-Mills equation for $A$ with $B=-*\,F_{A}$. Concerning Yang-Mills connections in two dimensions, Atiyah-Bott [Atiyah-Bott]*Theorem 6.7 gave a classification of principal bundles over Riemann surfaces that carry a Yang-Mills connection. So we ask whether Atiyah-Bott’s classification of Yang-Mills bundles in two dimensions can be extended to cone-flat bundles in $\dim M=2n\geq 2$ with respect to a closed, non-degenerate two-form $\zeta$. Indeed, we obtain the following for principal bundles which generalizes Atiyah-Bott’s result.
@@
As in the theorem, we will assume in the remainder of this section that $M$ is path-connected. We can give a more explicit description of $\Gamma$ in terms of path spaces. To do so, we first introduce some notations.
@@
which is a one-parameter family of subpaths, parametrized by $t$, all with the same starting point $\alpha(0)$, and ending point being at $\alpha(t)$.
The following subsets of $\Psi(M)$ will be of interest:
@@
    $\Psi^{a}_{a}(M)/\Psi^{a}_{a,0}(M)=\pi_{1}(M)$ is the fundamental group where the identity element is the equivalence class of the constant path at $a$, and the inverse reverses the path orientation, i.e. $\alpha^{-1}(s)=\alpha(1-s)\,$.
@@
    $\Psi^{a}_{a,0}(M)/\Psi^{a}_{a,\zeta}(M)\cong\mathbb{R}/{H}$. As $\zeta\neq 0$, for any real number $c$ there exists a contractible loop $\gamma$ that is the boundary of some disk $D$ such that $\int_{D}\zeta=c$. So $\Psi^{a}_{a,0}(M)/\Psi^{a}_{a,\zeta}(M)$ is equivalent to a quotient group $\mathbb{R}/{H}$ by the identification of a loop $\gamma\mapsto[\,\int_{D}\zeta\,]$, where
@@
    The group $H$ comes from noting that it is possible that the sum of two contractible curves $\partial D_{1}+\partial D_{2}$ can come from the vanishing boundary of a two-sphere formed by two hemisphere disks, $D_{1}$ and $D_{2}$, glued together at the equator such that $\partial D_{1}=-\partial D_{2}$.
@@
    $\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)=:\!\!\Gamma\,$. This is the extension of $\pi_{1}(M)$ by $\mathbb{R}/\overline{H}$ that appears in the statement of the theorem. In the case where $H$ is dense and thus $\mathbb{R}/H$ is not a Lie group, $\Gamma=\pi_{1}(M)$.
@@
Case 1. When $H^{+}$ is empty, $H=0$ and $\mathbb{R}/H=\mathbb{R}$. In this case, $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ is an $\mathbb{R}$-extension of $\pi_{1}(M)$. (For instance, this occurs for any Riemann surface $\Sigma_{g}$ with genus $g\geq 1$, since then $\pi_{2}(\Sigma_{g})=0\,$.)
Case 2. When $H^{+}$ has a minimal number, $H\simeq\mathbb{Z}$ and $\mathbb{R}/H\simeq S^{1}$. In this case $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ is an $S^{1}$-extension of $\pi_{1}(M)$. (This occurs for Riemann surface of genus $g=0$ and $\zeta$ is not an exact form.)
@@
We now proceed to prove the classification theorem Theorem 4.1. Our proof will be similar to that given by Morrison [Morrison] for Yang-Mills bundles in the two-dimensional case.
We will first consider the case that $\mathbb{R}/H$ is a Lie group. By the discussion above, $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ is either an $\mathbb{R}$ or an $S^{1}$ extension of $\pi_{1}(M)$.
@@
Step 1. Given a homomorphism $\rho:\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)\to G$, construct a corresponding principal $G$-bundle $P_{\rho}$ with a connection $A_{\rho}\,$.
Let $\Psi_{a}(M)$ be the space of classes of paths starting at $a\in M$, where any two paths $\delta_{1},\delta_{2}$ are identified if $\delta_{1}(1)=\delta_{2}(1)$ and $(\delta_{2})^{-1}\delta_{1}\in\Psi^{a}_{a,\zeta}(M)$. Notice that $\Psi_{a}(M)$ is a principal bundle over $M$. The projection to the base space $M$, $\tau:\Psi_{a}(M)\to M$, is given by $\delta\mapsto\delta(1)$, and its structure group is $\Gamma=\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$, since any element $\gamma\in\Psi^{a}_{a}(M)/\Psi^{a}_{a,\zeta}(M)$ can act on $\delta\in\Psi_{a}(M)$ on the right by $\delta\mapsto\delta\gamma$.
Given a homomorphism $\rho:\Gamma\to G$, we define the principal $G$-bundle $P_{\rho}$ as an associated $G$-bundle to $\Psi_{a}(M)$: